OpenStudy (anonymous):

Is the following always, sometimes, or never true? y − 6 = y

jimthompson5910 (jim_thompson5910):

Hint: y − 6 = y y − 6-y = y-y ... Subtract y from both sides. what's the next step?

OpenStudy (anonymous):

always

jimthompson5910 (jim_thompson5910):

what comes after y − 6-y = y-y

OpenStudy (anonymous):

y - 6= y

jimthompson5910 (jim_thompson5910):

that's going backwards

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

what is y-y

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

so y − 6-y = y-y turns into -6 = 0 is that true?

jimthompson5910 (jim_thompson5910):

is -6 = 0 true?

OpenStudy (anonymous):

no so is it jus y

jimthompson5910 (jim_thompson5910):

it's not true and it's never true so y − 6 = y is never true either